Interface of PetaSim and Emulators

March 3, 1998

nodeset.ini: (Nodeset description file)
name type number grainsize portcount (floatspeed) // CPU nodeset optional
node0_name portname0 link00_name portname1 link01_name ....
node1_name portname0 link10_name portname1 link11_name ....
....................
(total "number" of nodeset link pairs)
.............................
================================================
linkset.ini: (Linkset description file)
name type duplex number latency bandwidth
link0_name nodeset_00 nodeset_01
link1_name nodeset_!0 nodeset_11
.....................
(total "number" of nodeset-link-nodeset pairs)
..............................
================================================
dataset.ini: (Dataset description file)
name type totalsize bytesperunit floatperunit operationsperunit
................
=================================================

dist.ini: (Distribution description file)
name type
........

For examples:
nodeset.ini:
cpu CPU 4 64 1 1.0e-8
cpu0 cache0 link00
cpu1 cache0 link01
cpu2 cache1 link02
cpu3 cache1 link03
cache Cache 2 1048576 3
cache0 cpu0 link00 cpu1 link01 mem0 link10
cache1 cpu2 link02 cpu3 link03 mem1 link11
mem Memory 2 134217628 3
mem0 cache0 link10 diskctl0 link20 network link40
mem1 cache1 link11 diskctl1 link21 network link41
diskctl Disk 2 13421762800 4
diskctl0 mem0 link20 d0 link30 d1 link31 d2 link32
diskctl1 mem1 link21 d3 link33 d4 link34 d5 link35
disks Disk 6 13421762800 1
d0 diskctl0 link30
d1 diskctl0 link31
d2 diskctl0 link32
d3 diskctl1 link33
d4 diskctl1 link34
d5 diskctl1 link35
network Swith 1 0 2
network0 mem0 link40 mem1 link41

linkset.ini:

link0 Updown Full 4 0.001 104857600
link00 cpu0 cache0
link01 cpu1 cache0
link02 cpu2 cache1
link03 cpu3 cache1
link1 Updown Full 2 0.001 104857600
link10 cache0 mem0
link11 cache1 mem1
link2 Updown Full 2  0.05 10485760
link20 mem0 diskctl0
link21 mem1 diskctl1
link3 Updown Full 6 0.05 10485760
link30 diskctl0 d0
link31 diskctl0 d1
link32 diskctl0 d2
link33 diskctl1 d3
link34 diskctl1 d4
link35 diskctl1 d5
link4 Updown Full 2 0.1 512000
link40 mem0 network
link41 mem1 network

hfehrwf gjriejgre hjrieohjtj 4